OpenStudy (littlet):

Find the point that splits segment CD in half if point C is located at (−2, 4) and point D is located at (3, 7). (2.5, 1.5) (1, 5) (0.5, 5.5) (3, 2)

Mehek (mehek14):

use the midpoint formula: \((\dfrac{x_1+x_2}{2},\dfrac{y_1+y_2}{2})\)

Mehek (mehek14):

so for x we have -2 and 3 for y we have 4 and 7 \(\dfrac{-2+3}{2}=?\) \(\dfrac{4+7}{2}=?\)

OpenStudy (littlet):

-2+3= 1/2 = 0.5 4+7=11/2= 5.5

OpenStudy (hanamalik):

@lickking

OpenStudy (hanamalik):

@LickKing

Mehek (mehek14):

so what would be the answer then littlet?

OpenStudy (hanamalik):

* @LichKing

OpenStudy (littlet):

so would i multiply or divide?

Mehek (mehek14):

you already divided and got 0.5 and 5.5

OpenStudy (littlet):

Oh wait.. wow lol. answer C sorry lol